### v3

# FARM: Functional Group-Aware Representations for Small Molecules

Abstract
We introduce Functional Group-Aware Representations for Small Molecules (FARM), a novel foundation model designed to bridge the gap between SMILES, natural language, and molecular graphs. The key innovation of FARM lies in its functional group-aware tokenization, which directly incorporates functional group information into SMILES, enriching SMILES with detailed chemical context. For example, instead of using “O” to represent all oxygen atoms, we use specific tokens like “O_ketone” and “O_hydroxyl” to differentiate oxygen atoms belonging to distinct functional groups. This tokenization expands the chemical lexicon, effectively bridging the gap between SMILES and natural language in terms of vocabulary size, ultimately enhancing the model’s ability to predict molecular properties. FARM also represents molecules from two perspectives: by (1) using masked language modeling to capture atom-level features and (2) employing graph neural networks to encode the whole molecule topology. FARM leverages contrastive learning to aligns these two views of representations into a unified molecular embedding. We rigorously evaluate FARM on the MoleculeNet dataset, where it achieves state-of-the-art performance on 11 out of 13 tasks. These results highlight FARM’s potential to improve molecular representation learning and demonstrate its strong transfer learning capabilities, paving the way for promising applications in drug discovery and pharmaceutical research. The code is available at: https://github.com/thaonguyen217/farm_molecular_representation.
1 Introduction
Artificial intelligence (AI) has emerged as a transformative tool in accelerating scientific discovery, particularly in drug development. It is increasingly employed for tasks such as molecular property prediction, drug-target interaction prediction, and quantitative structure-activity relationship (QSAR) modeling (Chen et al., 2016; Wen et al., 2017; Shen and Nicolaou, 2019; Walters and Barzilay, 2020; Achary, 2020; Wang et al., 2022a; Edwards et al., 2022; Zhang et al., 2023a; Nguyen et al., 2024a; Edwards et al., 2024b, a). However, one of the central challenges in this field is the scarcity of large labeled datasets required for traditional supervised learning methods. This has shifted the focus towards self-supervised pre-trained models that can extract meaningful patterns from vast amounts of unlabeled molecular data (Shen and Nicolaou, 2019). As a result, the development of robust foundation models for molecular representations is now more critical than ever.
Despite significant advancements in other domains, such as natural language processing (NLP) and computer vision, there is still no dominant foundation model tailored to molecular representation in drug discovery (Zhang et al., 2023b). This paper begins to address this pressing gap by introducing an innovative approach that leverages functional group (FG)-aware tokenization in the context of both sequence-based and graph-based molecular representations. In the literature, terms such as “motifs,” “fragments,” “substructures,” and “building blocks” are often used interchangeably with “functional groups.” However, in this paper, we use the term “functional groups” to represent a subset of these molecular concepts, with a more rigorous definition grounded in chemical principles. A functional group refers to a chemically meaningful portion of a molecule that significantly influences its properties and behavior. This can include simple functional groups, such as hydroxyl (-OH), as well as more complex molecular substructures, such as ring systems, which serve distinct functional roles within the molecule. Functional groups (FGs) play a crucial role in determining a molecule’s properties. For instance, salicylic acid and aspirin share the same core structure but differ by just one functional group—a hydroxyl group in salicylic acid is replaced by an acetyl group in aspirin. This seemingly minor change results in substantial differences in their chemical properties and biological activities. The acetyl group in aspirin enables it to irreversibly inhibit cyclooxygenase (COX) enzymes, providing it with blood-thinning properties that make it effective in preventing heart attacks and strokes—functions that salicylic acid does not exhibit.
The innovation here lies in expanding tokenization granularity for SMILES and reducing tokenization granularity for molecular graphs in a way that explicitly interfaces with key drivers of molecular function. This approach enables molecular embeddings to more effectively support models in understanding and predicting molecular functions. For example, instead of representing all oxygen atoms with the same token (e.g., “O”), we use more specific tokens such as “O_ketone” and “O_hydroxyl” to distinguish oxygen atoms that belong to different functional groups in the SMILES string. This approach addresses a major limitation of sequence-based models, which typically focus only on individual atom types while neglecting the higher-level functional groups crucial for accurate molecular understanding. FG-aware tokenization enriches the SMILES representation with chemically relevant context, bridging the gap between the expansive vocabulary of natural language—the language for which Transformers were originally designed—and the limited chemical lexicon of the SMILES language, thereby reducing negative transfer during learning.
Conversely, instead of representing a ketone group (C=O) using two separate nodes in a graph, we represent it as a single node corresponding to the functional group as a whole. This abstraction yields a high-level “skeleton” of the molecule, where each node denotes a chemically meaningful functional unit rather than individual atoms. To capture interactions between these units, we construct a functional group (FG) knowledge graph that encodes chemically informed embeddings for each functional group based on their properties and co-occurrence patterns. These embeddings are then leveraged in a link prediction task to learn structural representations of molecules—allowing the model to infer both functional group similarity and likely connectivity. This FG graph captures structural relationships that are often implicit or obscured in SMILES representations. As a result, the functional group graph provides a complementary perspective to the FG-enhanced SMILES, encoding topological and structural context, while the enhanced SMILES preserves fine-grained sequence-level and chemical detail. Contrastive learning will then be used to align these two views, encouraging the model to learn representations that are sensitive to chemically meaningful structures. Figure 1 and Figure 2a present an overview of our molecular representation learning model.
Figure 1: FARM’s molecular representation learning framework. Given a molecule as input, we apply FG-aware tokenization to its SMILES representation and FG-based fragmentation to its molecular graph. Each resulting representation is then processed through a self-supervised learning objective. Finally, the two views are aligned using contrastive learning to make the final embedding robust by capturing both atom-level information from the SMILES and structural information from the graph.
In summary, our key contributions include:
  • FG-aware tokenization and fragmentation: We introduce FG-aware tokenization and fragmentation, adding rich chemical context to each atom and bridging the gap between SMILES and natural language.
  • Structural representation learning: We leverage a FG knowledge graph for robust FG embeddings and effectively learn molecular structure through link prediction between FGs.
  • Atom-feature and structural representation integration: We combine masked language model for learning atom-level features with GNNs for capturing structural information.
  • Robustness in downstream tasks: FARM demonstrates strong transfer learning capabilities - the core goal of pretrained models - outperforming other methods in 11 out of 13 tasks from the MoleculeNet benchmark.
2 Related Work
Functional group-Aware Molecular Representations   In recent years, there has been growing recognition that incorporating functional group (FG) information into molecular representations can significantly enhance model performance in downstream tasks. Approaches in this domain can be broadly classified into two categories: those leveraging language models (LMs) (Li et al., 2023; Xia et al., 2023) and those utilizing graph neural networks (GNNs) (Zhang et al., 2020, 2021; Yu and Gao, 2022; Yang et al., 2022; Wang et al., 2023; Wu et al., 2023; Han et al., 2023; Fang et al., 2023). Within these categories, methods either employ rule-based functional group detection, relying on predefined chemical rules to identify FGs, or adopt unsupervised strategies that infer substructures or motifs from the data. Regardless of the approach, these models consistently demonstrate that enriching molecular representations with functional group information leads to improved performance across a wide range of molecular property prediction tasks (Fang et al., 2023; Han et al., 2023; Wang et al., 2023). This underlines the importance of functional group awareness in achieving accurate and generalizable molecular representations.
In studies such as Zhang et al. (2021); Han et al. (2023); Chen et al. (2024); Yang et al. (2022), the BRICS algorithm (Degen et al., 2008) is employed to fragment molecules, with some extending the approach through additional rules to achieve finer-grained segmentation. However, this method focuses on general reaction-based bond breaking, rather than targeting functional group-specific characteristics. Additionally, while BRICS groups atoms into fragments, it does not explicitly identify or label functional groups, limiting its utility in tasks that require functional groups-specific information. Other works (Li et al., 2023; Chen et al., 2024) leverage RDKit (Landrum, 2010) for functional group detection. While RDKit is effective in identifying common functional groups, its capabilities are limited to well-known groups and do not extend to detecting more complex functional groups, such as ring systems that often dominate molecular datasets and play a critical role in molecular function. In this work, we present a rule-based functional group detection algorithm that accurately identifies 101 common groups and all ring-containing ones. Unlike frequent subgraph mining or motif-based methods, our approach is explicitly grounded in chemical principles, ensuring reliable and chemically meaningful detection aligned with expert definitions.
Contrastive Learning-Based Molecular Representations   Contrastive learning, a self-supervised learning technique, aims to learn representations by maximizing agreement between positive pairs while distinguishing them from negative samples. In the context of molecular representation, Wang et al. (2022b) applies this technique by augmenting each molecular graph to create slightly different versions, which are then treated as negative examples to enhance the learning process. In Pinheiro et al. (2022), the authors use a molecular graph encoder and a SMILES encoder to encode both molecular graphs and SMILES, employing contrastive loss to maximize the agreement between these embeddings. This method enriches SMILES with topology information and the molecular graph with sequence context, making the final embeddings of molecules more robust. In Zhang et al. (2020), the authors minimize the distance between the representation of a molecule and the representations of its constituent substructures. Collectively, these works highlight the versatility of contrastive learning in unifying diverse molecular representations, leading to improved downstream performance in molecular tasks. Our method extends this approach by using contrastive learning to maximize the agreement between FG-enhanced SMILES representations and FG graphs, thereby enhancing the alignment between sequence-based and graph-based representations, ensuring a more comprehensive and chemically informed embedding.
Figure 2: (a) FARM’s molecular representation learning model architecture. (b) Functional group-aware tokenization and fragmentation algorithm. (c) Snapshot of the functional group knowledge graph. (d) Generation of negative samples for contrastive learning.
3 Methodology
We present our methodologies for enhancing molecular representation learning. Section 3.1 introduces FG-aware tokenization and fragmentation, which injects detailed chemical context into both sequence and graph representations. Section 3.2 delves into the masked atom prediction task, a self-supervised technique that enables the model to learn atom-level representations. In Section 3.3, we describe our method for capturing the core molecular structure. Finally, Section 3.4 presents contrastive learning, which aligns FG-enhanced SMILES strings with FG graph embeddings to achieve a comprehensive, unified molecular representation.
3.1 Functional Group-Aware Tokenization and Fragmentation
We propose an FG-aware tokenization and fragmentation method that embeds detailed functional group information into molecular representations, tailoring it for both SMILES and graph-based models. This method defines a set of functional groups and employs an algorithm to detect them within the molecular graph. The algorithm traverses the graph, evaluating each atom based on criteria such as atom type, neighboring atom types and corresponding bonds, number of neighbors, atom charge, and bonded hydrogen atoms to identify the functional group. This approach ensures precise detection of functional groups within molecular graphs. For instance, a carbon atom with a charge of 0 and three neighbors, including a double-bonded oxygen atom, is classified as part of a ketone group (RCOR’), with the oxygen also contributing to the group. Additionally, we address cases where a functional group may be a subset of another. The algorithm first checks for the presence of the larger functional group. If it’s not identified, the algorithm then checks for smaller functional groups, ensuring correct identification even in complex structures. Appendix B.1 provides a full list of 101 non-ring-containing functional groups.
The results of the FG detection algorithm are utilized for two distinct processes:
  • FG-aware tokenization: The molecular graph, where each node (atom) is assigned to a specific functional group, is converted back into a SMILES string that incorporates functional group information. For example, a ketone-containing group originally depicted in SMILES as *C(=O)* is transformed into an FG-enhanced SMILES string like *C_ketone(=O_ketone)*. This FG-enhanced SMILES embeds additional chemical context directly into the molecular representation while remaining fully compliant with traditional SMILES rules (if the FG information is removed, the FG-enhanced SMILES reverts to its standard SMILES form). Similar to how words like “bank” have multiple meanings depending on context (e.g., river bank vs. financial bank), our method captures fine-grained semantic distinctions by representing functional groups with more precise tokens—for example, distinguishing ketone oxygen from hydroxyl oxygen.
  • FG-aware fragmentation: Once functional groups are identified, the molecule is segmented based on the bonds connecting these groups, as illustrated in Figure 2(b). These bonds are represented as edges in a graph, where the nodes correspond to the functional groups. This structure, known as the functional group graph, conveys the molecule’s structural information.
3.2 Atom-Level Representation Learning
We employ a masked language model architecture that takes FG-enhanced SMILES as input, using atom-level tokenization and leveraging masked atom prediction as a self-supervised task to train the model. Specifically, we adopt the BERT (Devlin, 2018) with the following loss:
$\mathcal{L}_{\text{MLM}}=-\sum_{i\in\mathcal{M}}\log P_{\theta}(x_{i}\mid x_{\setminus i}),\vskip-1.99997pt$
where $\mathcal{M}$ denotes the set of indices corresponding to the masked tokens, $x_{i}$ is the original token at position $i$, and $x_{\setminus i}$ represents the input sequence with the token at position $i$ masked. To evaluate the impact of masking, we conduct experiments with varying masking percentages and test the model’s performance on six downstream tasks. The results indicate that a masking percentage of 35% yields the highest average performance across tasks. Detailed results are provided in the Appendix D.1.
We examine the attention mechanism of the BERT model trained with FG-enhanced SMILES by visualizing the attention scores for a query atom (Figure 3). The attention map reveals that the model pays more attention to atoms that are strongly connected to the query atom than to those that are merely adjacent in the SMILES string. In detail, the query atom at position 23 shows higher attention to the atom at position 0, which is part of the same ring, rather than to the atom at position 26, which is closer in the SMILES string but not directly connected. This demonstrates that the model effectively learns the syntax and semantics of SMILES, capturing the underlying molecular structure rather than merely relying on the linear sequence of SMILES.
Figure 3: Visualization of the attention map of the BERT model trained with functional group-enhanced SMILES, demonstrating the model’s ability to learn and recognize the syntax of SMILES.
Figure 4: (a) Visualization of functional group knowledge graph embedding space: Clusters of five functional groups with closely related embeddings. (b) Link prediction performance: Substituting one functional group in a molecule with another generates parallel results across different molecules.
3.3 Molecular Structure Learning
To learn the structural view of molecules, we employ a two-step process. First, we derive FG embeddings from the FG knowledge graph. Next, we use these embeddings as inputs for a link prediction model to predict interactions between FGs. This approach effectively captures the relationships among functional groups, both in terms of their structural characteristics and interactions, thereby facilitating the implicit learning of molecular structure.
FG knowledge graph. The FG knowledge graph models each FG with various relations, such as the atoms, bonds, and ring structures it contains, as well as properties like water solubility, lipophilicity (logP), and more. A comprehensive list of FG knowledge graph relations is provided in the Appendix B.1. A snapshot of the FG knowledge graph is shown in Figure 2(c).
The FG knowledge graph embedding is learned by the ComplEx model (Trouillon et al., 2016) to obtain node embeddings. ComplEx is a matrix factorization model specifically designed to learn embeddings from multi-relational data. It is particularly effective at capturing complex, asymmetric relationships and handling one-to-many relations in knowledge graphs, making it well-suited for modeling intricate relational structures. Specifically, each element in a triple $(\mathbf{h},\mathbf{r},\mathbf{t})$ — where $\mathbf{h}$ is the head entity, $\mathbf{r}$ is the relation, and $\mathbf{t}$ is the tail entity — is represented as a complex vector. The score for a given triple $(\mathbf{h},\mathbf{r},\mathbf{t})$ is calculated as: $f(\mathbf{h},\mathbf{r},\mathbf{t})=\text{Re}\left(\mathbf{h}^{T}\mathbf{r}\cdot\mathbf{t}\right)$. ComplEx employs a margin-based ranking loss function defined as:
$\mathcal{L}_{\text{Graph}}=\sum_{\begin{subarray}{c}(\mathbf{h},\mathbf{r},\mathbf{t})\\ \in E^{+}\end{subarray}}\sum_{\begin{subarray}{c}(\mathbf{h^{\prime}},\mathbf{r},\mathbf{t^{\prime}})\\ \in E^{-}\end{subarray}}\max\left(0,\gamma+f(\mathbf{h^{\prime}},\mathbf{r},\mathbf{t^{\prime}})-f(\mathbf{h},\mathbf{r},\mathbf{t})\right)$ (1)
where $E^{+}$ denotes the set of positive triples, $E^{-}$ denotes the set of negative triples, and $\gamma$ represents the margin. Optimizing this loss function will minimize the score of positive triples $(\mathbf{h},\mathbf{r},\mathbf{t})$, while maximizing the score of negative triples $(\mathbf{h^{\prime}},\mathbf{r},\mathbf{t^{\prime}})$, with a margin $\gamma$ separating the two. This is achieved through margin-based ranking loss, where the function $f(\mathbf{h},\mathbf{r},\mathbf{t})$ evaluates the plausibility of the triples. The goal is to ensure that the score for positive triples is higher than that for negative triples by at least the margin $\gamma$, thus pushing positive triples closer and negative triples farther apart in the embedding space. The detailed implementation of ComplEx is described in Appendix D.2.
By embedding FGs through the knowledge graph, the model can capture both structural and property-based features of each FG, leading to richer FG representations. Figure 4(a) illustrates clusters of FGs in the FG knowledge graph embedding space, showing that similar FGs are closely positioned, suggesting effective structural and property-based grouping.
Link prediction. For link prediction with a graph convolutional network (GCN), we first segment molecules into functional groups through FG-aware molecular segmentation, connecting each group via single bonds. We then utilize embeddings from the FG knowledge graph as node features for the GCN. The training involves computing node embeddings through graph convolution:
$\mathbf{h}_{i}^{\prime}=\text{ReLU}\left(\mathbf{W}\cdot\frac{1}{|\mathcal{N}(i)|}\sum_{j\in\mathcal{N}(i)}\mathbf{h}_{j}\right)$
where $\mathbf{h}_{i}^{\prime}$ is the updated embedding for node $i$, computed by averaging the embeddings $\mathbf{h}_{j}$ of neighboring nodes $\mathcal{N}(i)$, applying the weight matrix $\mathbf{W}$, and passing through ReLU activation function. The score estimates the probability of connections between nodes is computed with a multi-layer perceptron (MLP) and a sigmoid function: $p_{ij}=\sigma(\text{MLP}(\mathbf{h}_{i}\oplus\mathbf{h}_{j}))$.
We then sample positive edges $E^{+}$ and negative edges $E^{-}$, optimizing the model to maximize scores for positive edges and minimize those for negative ones using the loss function:
$\mathcal{L}_{\text{Link}}=-\frac{1}{|E^{+}|}\sum_{(i,j)\in E^{+}}\log p_{ij}-\frac{1}{|E^{-}|}\sum_{(i,j)\in E^{-}}\log(1-p_{ij})$
Figure 4(b) and Figure 8 (Appendix D.3) demonstrates the capability of our molecular structure representation model, showing that, akin to word pair analogy tasks in NLP, replacing one functional group in a molecule with another (in this case, replacing -OH with -COOH) produces parallel results across different molecules. This demonstrates the model’s ability to effectively capture and preserve chemical analogies, highlighting its robustness in learning and representing molecular structures.
3.4 Molecular Structure Integration via Contrastive Learning
To integrate FG-enhanced SMILES representations with molecular structure information, we employ contrastive learning to align these representations with FG graph embeddings. This method captures both the atom-level and topological aspects of molecular structures, allowing the model to develop a unified representation that integrates chemical context with overall molecular architecture.
In this framework, each molecule is treated as a pair of representations: the FG-enhanced SMILES and its corresponding FG graph. The contrastive learning task encourages the embeddings of these two representations to be as similar as possible for the same molecule, while pushing apart the representations of different molecules. This allows the model to capture both local chemical features (from the FG-enhanced SMILES) and global molecular topology (from the FG graph).
To enhance the learning process and make it more robust, we generate negative examples by augmenting the FG graph. We apply two types of augmentations: (1) node deletion, where one or more functional groups are removed from the graph, and (2) node swapping, where functional groups are randomly exchanged with one another. Figure 2(d) illustrates how these augmentations are applied to generate negative examples from a FG graph. These augmentations create harder negative examples that force the model to better understand the correct structure and connectivity between functional groups, making it more effective at learning meaningful molecular representations.
Specifically, given a positive pair $(\mathbf{h}_{\text{MLM}},\mathbf{h}_{\text{pos}})$, where $\mathbf{h}_{\text{MLM}}$ is the atom-level representation derived from a pretrained BERT model and $\mathbf{h}_{\text{pos}}$ is the corresponding structure representation from a graph neural network (GNN), and a negative pair $(\mathbf{h}_{\text{MLM}},\mathbf{h}_{\text{neg}})$, where $\mathbf{h}_{\text{neg}}$ is a augmented FG-graph, the contrastive loss can be written as: $\mathcal{L}_{\text{CL}}=\frac{1}{N}\sum_{i=1}^{N}\max\left(0,\gamma-\text{cos}(\mathbf{h}_{\text{MLM}},\mathbf{h}_{\text{pos}})+\text{cos}(\mathbf{h}_{\text{MLM}},\mathbf{h}_{\text{neg}})\right)$.
where $\gamma$ is the margin parameter and $N$ is the number of training examples (or contrastive pairs). The final objective function for integration is: $\mathcal{L}_{\text{Integration}}=\lambda_{\text{MLM}}\cdot\mathcal{L}_{\text{MLM}}+\lambda_{\text{CL}}\cdot\mathcal{L}_{\text{CL}}$, where $\lambda_{\text{MLM}}$ and $\lambda_{\text{CL}}$ are hyperparameters that control the contribution of each loss to the overall objective. Details on the implementation of contrastive training can be found in the Appendix D.4.
4 Pre-training Data Collection and Diversity Assessment
The performance and generalizability of a machine learning model is heavily dependent on the quality and diversity of its training data. Given the vastness of chemical space, it is vital that the training data represents a sufficiently representative subset of this space. Traditional approaches assess the diversity of datasets based on criteria such as the number of chemical elements and the number of atoms per molecule. However, to the best of our knowledge, no previous studies have systematically evaluated the diversity of large chemical datasets like ZINC (Irwin and Shoichet, 2005) or ChEMBL (Gaulton et al., 2011) in terms of functional groups. This gap is significant, as functional groups provide deep insights into the structural and functional complexity of molecules.
In this work, we introduce a novel criterion to assess dataset diversity by using the size of the FG-enhanced SMILES lexicon as a metric. By analyzing the size of this lexicon, we can assess whether the dataset captures a comprehensive range of chemical functionalities, which is crucial for building robust molecular foundation models. Table 1 presents the size of the FG-enhanced SMILES lexicon for the ZINC15 and ChEMBL25 datasets, as well as the dataset we collected for training our foundation model. The table shows that despite its widespread use for training foundation models in small molecule representation, ZINC15 is significantly less diverse compared to ChEMBL25. This is largely due to the fact that ZINC15 is primarily designed to include molecules that adhere to Lipinski’s Rule of Five for drug-likeness, excluding more exotic or less common elements that are less relevant to pharmaceutical chemistry. This limited diversity may negatively impact the model’s performance on out-of-distribution datasets, which contain a broader range of atom types and functional groups not present in ZINC15. In contrast, ChEMBL25 is far more diverse and thus better suited for training foundation models that can be fine-tuned for various downstream tasks. Based on this insight, we collected a dataset that includes the entire ChEMBL25 database, libraries from chemical drug suppliers, and a subset of ZINC15 to ensure a more comprehensive coverage of chemical space. Detailed information about the collected data can be found in the Appendix A.
Table 1: Evaluation of the FG-enhanced SMILES lexicon size across various databases.
[…]
ZINC15 3M 10 [5; 63] 1,151 1,089
ChEMBL25 1.8M 35 [1; 867] 15,269 10,016
Collected dataset 20M 46 [1; 867] 22,364 14,741
Table 2: Performance comparison of FARM and baseline models on MoleculeNet classification tasks. The first 10 models incorporate FG (functional group) information to enhance representation learning.
Physiology Biophysics
Dataset BBBP Tox21 ToxCast SIDER ClinTox BACE MUV HIV
#tasks 1 12 617 27 2 1 17 1
#samples 2,039 7,831 8,575 1,427 1,478 1,513 93,807 41,127
Evaluation Metric ROC-AUC (%) ($\uparrow$) ROC-AUC (%) ($\uparrow$)
MICRO (Zhang et al., 2020) 84.4 $\pm$ 1.1 77.0 $\pm$ 0.8 65.2 $\pm$ 0.8 56.7 $\pm$ 0.9 77.0 $\pm$ 2.0 77.2 $\pm$ 2.0 - 75.1 $\pm$ 1.1
MGSSL (Zhang et al., 2021) 69.7 $\pm$ 0.9 76.5 $\pm$ 0.3 64.1 $\pm$ 0.7 61.8 $\pm$ 0.8 80.7 $\pm$ 2.1 79.1 $\pm$ 0.9 78.7 $\pm$ 1.5 78.8 $\pm$ 1.2
MoleOOD (Yang et al., 2022) 71.0 $\pm$ 0.8 - - 63.4 $\pm$ 0.7 - 84.3 $\pm$ 1.1 - 79.4 $\pm$ 0.5
MCM (Wang et al., 2023) 90.0 $\pm$ 3.1 80.2 $\pm$ 1.5 - 62.7 $\pm$ 2.8 65.5 $\pm$ 1.4 82.0 $\pm$ 5.5 - -
HimGNN (Han et al., 2023) 92.8 $\pm$ 2.7 80.7 $\pm$ 1.7 - 64.2 $\pm$ 2.3 91.7 $\pm$ 3.0 85.6 $\pm$ 3.4 - -
FG-BERT (Li et al., 2023) 70.2 $\pm$ 0.9 78.4 $\pm$ 0.8 63.3 $\pm$ 0.8 64.0 $\pm$ 0.7 83.2 $\pm$ 1.6 84.5 $\pm$ 1.5 75.3 $\pm$ 2.4 77.4 $\pm$ 1.0
HiMol (Zang et al., 2023) 71.3 $\pm$ 0.6 76.0 $\pm$ 0.2 - 62.5 $\pm$ 0.3 70.6 $\pm$ 2.1 84.6 $\pm$ 0.2 - -
Mole-BERT (Xia et al., 2023) 71.9 $\pm$ 1.6 76.8 $\pm$ 0.5 62.8 $\pm$ 1.1 62.8 $\pm$ 1.1 78.9 $\pm$ 3.0 80.8 $\pm$ 1.4 78.6 $\pm$ 1.8 78.2 $\pm$ 0.8
MolCLR (Wang et al., 2022b) 73.3 $\pm$ 1.0 74.1 $\pm$ 5.3 - 61.2 $\pm$ 3.6 89.8 $\pm$ 2.7 82.8 $\pm$ 0.7 78.9 $\pm$ 2.3 77.4 $\pm$ 0.6
GLAD (Nguyen et al., 2024b) 80.4 $\pm$ 1.5 - - 64.7 $\pm$ 1.8 87.3 $\pm$ 1.2 85.7 $\pm$ 0.9 - -
N-GRAM (Hu et al., 2019) 70.8 $\pm$ 1.5 78.7 $\pm$ 0.4 66.5 $\pm$ 0.3 62.7 $\pm$ 0.8 72.6 $\pm$ 1.5 84.5 $\pm$ 0.7 81.3 $\pm$ 2.1 79.9 $\pm$ 0.7
GROVER (Rong et al., 2020) 86.8 $\pm$ 2.2 80.3 $\pm$ 2.0 65.3 $\pm$ 0.5 61.2 $\pm$ 2.5 70.3 $\pm$ 13.7 82.4 $\pm$ 3.6 67.3 $\pm$ 1.8 68.2 $\pm$ 1.1
GraphMVP (Liu et al., 2021) 72.4 $\pm$ 1.6 75.9 $\pm$ 0.5 63.1 $\pm$ 0.4 63.1 $\pm$ 0.4 79.1 $\pm$ 2.8 81.2 $\pm$ 0.9 77.7 $\pm$ 0.6 77.0 $\pm$ 1.2
GEM (Fang et al., 2022) 88.8 $\pm$ 0.4 78.1 $\pm$ 0.4 69.2 $\pm$ 0.4 63.2 $\pm$ 1.5 90.3 $\pm$ 0.7 87.9 $\pm$ 1.1 75.3 $\pm$ 1.5 81.3 $\pm$ 0.3
UniMol (Zhou et al., 2023) 72.9 $\pm$ 0.6 79.6 $\pm$ 0.5 69.6 $\pm$ 0.1 65.9 $\pm$ 1.3 91.9 $\pm$ 1.8 85.7 $\pm$ 0.2 82.1 $\pm$ 1.3 82.8 $\pm$ 0.3
FARM (Ours) 93.3 $\pm$ 0.2 80.8 $\pm$ 1.1 69.9 $\pm$ 0.5 65.9 $\pm$ 0.7 82.2 $\pm$ 0.7 89.6 $\pm$ 0.4 82.7 $\pm$ 2.1 83.5 $\pm$ 0.5
Table 3: Performance comparison of FARM and baseline models on MoleculeNet regression tasks. Performance is evaluated by RMSE and MAE.
Physical chemistry Quantum mechanics
Dataset ESOL Freesolv Lipophilicity QM8 QM9
#tasks 1 1 1 12 3
#samples 1,128 642 4,200 21,786 133,885
Evaluation Metric RMSE ($\downarrow$) MAE ($\downarrow$)
HimGNN (Han et al., 2023) 0.870 $\pm$ 0.154 1.921 $\pm$ 0.474 0.632 $\pm$ 0.016 - -
FG-BERT (Li et al., 2023) 0.944 $\pm$ 0.025 - 0.655 $\pm$ 0.009 - -
Mole-BERT (Xia et al., 2023) 1.015 $\pm$ 0.003 - 0.676 $\pm$ 0.002 - -
N-GRAM (Hu et al., 2019) 1.100 $\pm$ 0.030 2.510 $\pm$ 0.191 0.880 $\pm$ 0.121 0.0320 $\pm$ 0.003 0.00964 $\pm$ 0.00031
GROVER (Rong et al., 2020) 1.423 $\pm$ 0.288 2.947 $\pm$ 0.615 0.823 $\pm$ 0.010 0.0182 $\pm$ 0.001 0.00719 $\pm$ 0.00208
GEM (Fang et al., 2022) 0.813 $\pm$ 0.028 1.748 $\pm$ 0.114 0.674 $\pm$ 0.022 0.0163 $\pm$ 0.001 0.00562 $\pm$ 0.00007
MolCLR (Wang et al., 2022b) 1.113 $\pm$ 0.023 2.301 $\pm$ 0.247 0.789 $\pm$ 0.009 0.0185 $\pm$ 0.013 0.00480 $\pm$ 0.00003
UniMol (Zhou et al., 2023) 0.788 $\pm$ 0.029 1.480 $\pm$ 0.048 0.603 $\pm$ 0.010 0.0156 $\pm$ 0.001 0.00467 $\pm$ 0.00004
FARM (Ours) 0.761 $\pm$ 0.031 1.097 $\pm$ 0.033 0.778 $\pm$ 0.005 0.0146 $\pm$ 0.001 0.00456 $\pm$ 0.00001
5 Experimental Results
Data and Splits   We consider 13 benchmark tasks in the MoleculeNet dataset (Wu et al., 2018). Following previous works, the data is split using a scaffold split into training, validation, and test sets with an 8:1:1 ratio, ensuring fair and consistent evaluation across all models.
Evaluation Metrics.   We use ROC-AUC as the evaluation metric for classification tasks due to the high imbalance in some datasets. For physical chemistry tasks (ESOL, Freesolv, and Lipophilicity), we use RMSE, and for quantum mechanics tasks (QM8 and QM9), we use MAE, following previous works. For each downstream task, we split the data using three random seeds, train the models, and report the average and standard deviation of the results.
Baselines   We consider works that incorporate functional group information to enhance representation learning (Zhang et al., 2020, 2021; Yang et al., 2022; Wang et al., 2023; Han et al., 2023; Li et al., 2023; Zang et al., 2023; Wang et al., 2022b; Xia et al., 2023; Nguyen et al., 2024b), and other approaches that utilize masked atom prediction as a self-supervised training task (Rong et al., 2020; Hu et al., 2019; Liu et al., 2021; Fang et al., 2022; Zhou et al., 2023).
Results   Table 2 and 3 present the performance of FARM alongside other baseline models on seven MoleculeNet classification tasks. FARM consistently outperforms baseline models on various benchmarks, demonstrating its robustness and versatility. Its strong performance across various classification and regression tasks indicates that FARM is a highly effective pretrained model, well-suited for a broad range of downstream tasks in molecular property prediction.
Ablation Study on FARM Component Analysis   To assess the effectiveness of each component in our architecture, we conducted a comprehensive ablation study across several MoleculeNet benchmark tasks. The results are shown in Table 10. For more details about the baselines, please refer to appendix F. The results show that each component of FARM contributes to improved performance compared to the baseline. Combining all components leads to strong overall performance of FARM across all test tasks.
Ablation Study on Fragmentation Methods   We compare our FG-based fragmentation method with BRICS fragmentation. When treating each BRICS fragment as a functional group, the enriched SMILES vocabulary size reaches 575,612 (min freq $\leq$ 5), compared to just 14,741 with our FG-based method. This is because BRICS produces larger, less reusable fragments, resulting in a bloated vocabulary with many rare tokens—making training harder and increasing the risk of out-of-vocabulary issues. As shown in Table 4, our FG-based method consistently outperforms BRICS across all six MoleculeNet tasks under the same setup.
Table 4: Performance comparison of the model using BRICS fragmentation versus our FG-based fragmentation across six MoleculeNet tasks.
BBBP BACE HIV Average ESOL FreeSolv Average QM9
#tasks 1 1 1 1 1 3
#samples 2039 1513 41127 1128 642 133885
Metric ROC-AUC ($\uparrow$) RMSE ($\downarrow$) MAE ($\downarrow$)
BRICS 88.87 $\pm$ 1.5 89.54 $\pm$ 1.09 70.09 $\pm$ 2.24 82.83 1.340 $\pm$ 0.235 0.823 $\pm$ 0.14 1.081 0.0085 $\pm$ 0.00034
OURS 96.23 $\pm$ 0.7 96.19 $\pm$ 0.65 82.13 $\pm$ 1.10 91.43 0.734 $\pm$ 0.039 0.308 $\pm$ 0.08 0.521 0.0038 $\pm$ 0.00014
6 Conclusion, Limitations, and Future Work
In summary, FARM demonstrates robust performance across various MoleculeNet classification tasks, outperforming or matching baseline models. The integration of functional group information and the alignment of FG-enhanced SMILES representations with FG graph embeddings through contrastive learning significantly enhance its effectiveness. While FARM shows strong performance, there is a key limitation that should be addressed in future work. The current model does not incorporate a full 3D molecular representation, which is critical for capturing stereochemistry and spatial configurations that influence molecular properties. Integrating 3D structural information Yan et al. (2024) could further improve the model’s ability to make accurate predictions.
Looking ahead, our ultimate goal is to develop a pre-trained atom embedding that parallels the capabilities of pre-trained word embeddings in NLP. This would enable a richer and more nuanced understanding of molecular properties and behaviors at the atomic level. Similarly, we aim to achieve molecule-level representations that are as expressive and versatile as sentence-level embeddings in NLP, capturing both local and global molecular features. By bridging the gap between atom-wise embeddings and holistic molecule representations, FARM paves the way for more accurate, generalizable molecular predictions across a variety of tasks.
Acknowledgements
This work is supported by the Molecule Maker Lab Institute: an AI research institute program supported by NSF under award No. 2019897. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied of, the National Science Foundation, and the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.
Appendix A Molecular Datasets
A.1 Training data
We collected a diverse dataset to train our FARM model from various sources, including ChEMBL25, ZINC15, and several chemical suppliers. The number of compounds in each dataset is reported as follows:
Table 5: List of compound suppliers and number of compounds
Supplier Number of Compounds Source
Targetmol 22,555 https://www.targetmol.com/
Chemdiv 1,741,620 https://www.chemdiv.com/
Enamine 862,698 https://enamine.net/
Life Chemical 347,657 https://lifechemicals.com/
Chembridge 1,405,499 https://chembridge.com/
Vitas-M 1,430,135 https://vitasmlab.biz/
InterBioScreen 560,564 https://www.ibscreen.com/
Maybridge 97,367 https://chembridge.com/
Asinex 601,936 https://www.asinex.com/
Eximed 61,281 https://eximedlab.com/
Princeton BioMolecular 1,647,078 https://princetonbio.com/
Otava 9,203,151 https://www.otava.com/
Alinda Chemical 733,152 https://www.alinda.ru/synthes_en.html
ChEMBL 25 1,785,415 https://www.ebi.ac.uk/chembl/
ZINC15 4,000,000 https://zinc15.docking.org/
Total 20,000,000
A.2 Lexicon Growth for FG-enhanced SMILES Dataset
As part of our investigation into the relationship between corpus size and vocabulary size, we conducted experiments to observe how vocabulary growth evolves as the dataset increases. Our results show that while the vocabulary size increases with the corpus size, it eventually stabilizes. Specifically, when expanding the corpus from 20 million tokens to 200 million tokens, the vocabulary grew from 14,741 to 42,364 tokens, and with a 1 billion-token corpus, the vocabulary reached a plateau of 42,391 tokens. This indicates that the lexicon tends to converge around 42.3K tokens as the corpus size increases, suggesting that further growth in corpus size may not result in substantial increases in vocabulary.
Even with the smaller vocabulary of 14,741 tokens derived from the 20 million-token dataset, the model was able to achieve excellent coverage of common tokens. In fact, the 14,741 tokens from the 20M dataset accounted for 99.997% of all tokens in the 200M dataset, with 2,932,883,078 out of 2,932,984,142 tokens already represented. This demonstrates that a relatively smaller vocabulary can effectively capture the majority of the data, even with larger corpora. Moreover, the ratio of the UNKNOWN token in downstream datasets remains consistently below 0.1
A.3 Downstream Tasks Data
In Table 6, we provide an overview of the datasets used for evaluating the performance of our model on various downstream tasks. Each dataset is denoted by its name, followed by the number of tasks it encompasses, the total number of samples available in each dataset, and a brief description. These datasets cover a range of chemical and biological properties, enabling comprehensive evaluation of the model’s performance across different tasks in molecular representation learning.
Table 6: Overview of downstream tasks, corresponding sample sizes, and dataset descriptions.
Dataset # Tasks # Samples Description
BBBP 1 2,039 Benchmark for Blood-Brain Barrier permeability prediction,
assessing whether compounds can cross the blood-brain barrier.
Tox21 12 7,831 Toxicology data containing multiple assays for evaluating
the toxicity of compounds across various endpoints.
SIDER 27 1,427 Side Effect Resource dataset that includes drug side effects
associated with FDA-approved drugs, focusing on adverse drug reactions.
ClinTox 2 1,478 Clinical Toxicology dataset designed to predict the toxicity of
drug-like compounds based on clinical data.
BACE 1 1,513 Data for predicting activity against the beta-secretase enzyme,
relevant for Alzheimer’s disease drug discovery.
MUV 17 93,807 Multiple Unrelated Variables dataset aimed at assessing the ability
to predict various molecular properties and activities.
HIV 1 41,127 Dataset focused on predicting the activity of compounds
against the HIV virus, crucial for antiviral drug development.
ESOL 1 1,128 Dataset used for estimating the solubility of organic compounds
in water, useful for understanding compound behavior in biological systems.
FreeSolv 1 642 Dataset containing free energy of solvation values for
small organic molecules in water, aiding in solvation energy predictions.
Lipophilicity 1 4,200 Data focused on predicting the octanol-water
partition coefficient, a key measure of a compound’s lipophilicity.
QM8 12 21,786 Quantum Mechanics dataset that provides a range of molecular properties
computed using quantum mechanical methods for small organic molecules.
QM9 3 133,885 Quantum Mechanics dataset providing molecular properties for
a large set of small organic compounds.
Appendix B FG-aware Tokenization and Fragmentation
B.1 The List of Functional Groups
The exhaustive list of 101 functional groups that can be detected by the functional group detection algorithm includes: Tertiary carbon, Quaternary carbon, Alkene carbon, Cyanate, Isocyanate, Hydroxyl, Ether, Hydroperoxy, Peroxy, Haloformyl, Aldehyde, Ketone, Carboxylate, Carboxyl, Ester, Hemiacetal, Acetal, Hemiketal, Ketal, Orthoester, Carbonate ester, Orthocarbonate ester, Amidine, Carbamate, Isothiocyanate, Thioketone, Thial, Carbothioic S-acid, Carbothioic O-acid, Thiolester, Thionoester, Carbodithioic acid, Carbodithio, Trifluoromethyl, Difluorochloromethyl, Bromodifluoromethyl, Trichloromethyl, Bromodichloromethyl, Tribromomethyl, Dibromofluoromethyl, Triiodomethyl, Difluoromethyl, Fluorochloromethyl, Dichloromethyl, Chlorobromomethyl, Chloroiodomethyl, Dibromomethyl, Bromoiodomethyl, Diiodomethyl, Alkyl, Alkene, Alkyne, Carboxylic anhydride, Primary amine, Secondary amine, Amide, Imide, Tertiary amine, 4-ammonium ion, Hydrazone, Primary ketimine, Primary aldimine, Secondary ketimine, Secondary aldimine, Nitrile, Azide, Azo, Nitrate, Isonitrile, Nitrosooxy, Nitro, Nitroso, Aldoxime, Ketoxime, Sulfhydryl, Sulfide, Disulfide, Sulfinyl, Sulfonyl, Sulfur dioxide, Sulfuric acid, Sulfino, Sulfonic acid, Sulfonate ester, Thiocyanate, Phosphino, Phosphono, Phosphate, Phosphodiester, Phosphoryl, Borono, Boronate, Borino, Borinate, Silyl ether, Dichlorosilane, Trimethylsilyl, Fluoro, Chloro, Bromo, Iod.
B.2 Naming functional groups with Rings in Fused Ring Systems
Fused ring systems are a diverse and prevalent class of functional groups, accounting for 99.37% of the total functional groups in our dataset (147,564 out of 148,507 FGs). Despite their importance, many of these systems lack standardized nomenclature. To address this, we propose a systematic approach to naming these ring systems based on their ring sizes and core structures.
Each ring in a fused ring system is named according to its size. For instance, a six-membered aromatic ring like benzene is named ring_6. This straightforward approach provides a clear identifier for individual rings within a system. For systems composed of multiple fused rings, we use the following naming convention:
  • Identification: Determine the smallest atom index for each ring within the system.
  • Sorting: Arrange the rings by increasing atom indices.
  • Construction: Combine the ring sizes in ascending order. For example, a fused system with a six-membered ring and a five-membered ring would be named ring_5_6.
This systematic naming helps in identifying and categorizing complex fused ring systems by focusing on their core structure. The core structure is defined as the central framework of interconnected rings that forms the fundamental backbone of the molecule. The core structure of a ring system is important because it influences the molecule’s reactivity, stability, and biological activity. In SMILES notation, which uses lowercase characters to indicate atoms within aromatic rings, we can enhance the representation by combining the atom symbol (uppercase or lowercase) with the core structure, thereby providing a comprehensive depiction of the ring system. Figure 5a illustrates an example of naming a fused ring system based on the rules described above, and Figure 5b shows how FG-aware tokenization is applied.
Figure 5: (a) Example of naming a fused ring system in 4 steps: generate the core structure of the functional group, index atoms using RDKit, select the smallest-index atom in each ring and sort, and name the fused ring system based on ring size. (b) Example of FG-aware tokenization.
After completing the naming process, we derive a new FG-enhanced SMILES representation for the molecules. We then analyze our collected dataset, which comprises 20 million samples of FG-enhanced SMILES, to evaluate the results. This dataset includes representations of 46 different elements. Notably, 11 elements are represented by only a single form, indicating their rare occurrence within the dataset (excluding hydrogen). These elements are: H, Ti, V, Cr, Rb, Mo, Rh, Sb, Ba, Pb, and Bi. In contrast, the remaining 35 elements feature at least two representations, each corresponding to distinct FGs. The distribution of these elements is visualized in Figure 6, highlighting the diversity of representations in our dataset. The most prevalent element in our dataset is Carbon, with 9,112 FGs containing it. Nitrogen follows as the second most prevalent element, represented in 2,549 FGs, while Oxygen and Sulfur appear in 2,156 and 571 FGs, respectively.
Figure 6: Number of functional groups associated with different chemical elements in the FG-enhanced SMILES dataset. The y-axis represents the natural logarithm (log, base $e$) of the count.
Appendix C FG Knowledge Graph
The FG knowledge graph is designed to capture both the structural and property-related information of FGs. The list of relations includes:
Table 7: Key relations defined in the FG knowledge graph. (Note: Continuous values, such as LogP and water solubility, are discretized by rounding to the nearest integer.)
Relation Description
contain_atom Identifies atoms present in the FG (e.g., C, H, O, N).
contain_bond Specifies types of bonds in the FG (e.g., single, double, triple, aromatic).
functional_group Recognizes functional groups in the FG (e.g., hydroxyl, carboxyl, amine).
contain_ring_[n] Indicates the presence of a non-aromatic ring of size n in the FG.
contain_aromatic_ring_[n] Indicates the presence of an aromatic ring of size n in the FG.
num_substitutes Specifies the number of substituents (e.g., alkyl or aryl groups) in the FG.
is_hydrogen_bond_donor Identifies whether the FG contains a functional group capable of donating hydrogen bonds.
is_hydrogen_bond_acceptor Identifies whether the FG contains a functional group capable of accepting hydrogen bonds.
logp Measures the lipophilicity of the FG using the logP value (calculated via RDKit).
In the collected dataset, values range from -35 to 31.
water_solubility Predicts the solubility of the FG in water, based on logP, molecular weight, and TPSA.
In the collected dataset, values range from -5 to 8.
core_smiles The SMILES representation of the core structure of the FG.
  • List of functional groups that act as hydrogen bond donors: Hydroxyl, Hydroperoxy, Primary amine, Secondary amine, Hydrazone, Primary ketimine, Secondary ketimine, Primary aldimine, Amide, Sulfhydryl, Sulfonic acid, Thiolester, Hemiacetal, Hemiketal, Carboxyl, Aldoxime, Ketoxim.
  • List of functional groups that act as hydrogen bond acceptors: Ether, Peroxy, Haloformyl, Ketone, Aldehyde, Carboxylate, Carboxyl, Ester, Ketal, Carbonate ester, Carboxylic anhydride, Primary amine, Secondary amine, Tertiary amine, 4-Ammonium ion, Hydrazone, Primary ketimine, Secondary ketimine, Primary aldimine, Amide, Sulfhydryl, Sulfonic acid, Thiolester, Aldoxime, Ketoxi.
Appendix D Implementation Details
D.1 Training Masked Language Model for SMILES Representation
We trained the BERT model using Hugging Face [Wolf et al., 2020] on the masked molecule prediction task with both conventional SMILES and FG-enhanced SMILES from our collected dataset. To assess the impact of different masking percentages, we trained BERT models with masking percentages of 0.15, 0.25, 0.35, 0.45, and 0.55. The models were then evaluated on seven MoleculeNet tasks, including three classification tasks and four regression tasks, to determine the optimal masking percentage. The results, presented in Table 8, indicate that a masking percentage of 0.35 yields the best performance across the considered downstream tasks.
Table 8: Performance of BERT models with varying masking percentages across six MoleculeNet tasks. The data is split using a random split into training, validation, and test sets with an 8:1:1 ratio.
BBBP BACE HIV Average ESOL FreeSolv Average QM9
#tasks 1 1 1 1 1 3
#samples 2039 1513 41127 1128 642 133885
Metric ROC-AUC ($\uparrow$) RMSE ($\downarrow$) MAE ($\downarrow$)
0.25 93.01 $\pm$ 0.9 94.31 $\pm$ 1.08 80.17 $\pm$ 1.5 89.16 0.688 $\pm$ 0.033 0.622 $\pm$ 0.007 0.655 0.0091 $\pm$ 0.00001
0.25 93.59 $\pm$ 1.7 93.94 $\pm$ 1.4 81.03 $\pm$ 1.9 89.52 0.543 $\pm$ 0.030 0.714 $\pm$ 0.010 0.629 0.0032 $\pm$ 0.00001
0.35 94.36 $\pm$ 0.5 94.54 $\pm$ 0.4 81.93 $\pm$ 1.7 90.27 0.608 $\pm$ 0.031 0.507 $\pm$ 0.030 0.558 0.0041 $\pm$ 0.00001
0.45 93.48 $\pm$ 1.3 94.36 $\pm$ 0.90 80.12 $\pm$ 1.7 89.32 0.795 $\pm$ 0.028 0.493 $\pm$ 0.008 0.644 0.0048 $\pm$ 0.00001
0.55 92.85 $\pm$ 1.1 88.68 $\pm$ 1.0 79.89 $\pm$ 0.90 87.14 0.734 $\pm$ 0.030 0.599 $\pm$ 0.005 0.667 0.0097 $\pm$ 0.00001
Additional details of the training setup include training the BERT model on 20 million SMILES for 15 epochs using two NVIDIA Tesla V100 GPUs. The learning rate was set to $1e-5$ , with a batch size of 128, and model checkpoints were saved after every 10,000 batches. This setup was also applied to the baseline model, which used conventional SMILES for comparison.
Figure 7 illustrates the convergence behavior of the models trained on different representations of molecular data. The model utilizing FG-enhanced SMILES exhibits a slower convergence rate, attributed to the increased complexity of its vocabulary, reflecting its closer resemblance to natural language. The SMILES model converges by step 200 (after processing 25,600 SMILES), while the FG-enhanced SMILES model achieves convergence by step 300 (after processing 38,400 SMILES). Notably, despite the larger prediction vocabulary (14,714 vs. 93), the FG-enhanced model ultimately reaches a lower loss, suggesting its enhanced capacity to capture intricate molecular representations and improve generalization in complex tasks. This indicates the model’s ability to leverage functional group information effectively, potentially leading to better performance in downstream applications.
Figure 7: Loss curves for the masked language model (MLM) during training on two datasets: standard SMILES and functional group-enhanced SMILES.
D.2 Training FG Knowledge Graph Embedding Model for Molecular Structure Representation
Once the FG knowledge graph is constructed as detailed in Section C, we utilize the ComplEx model to learn embeddings for the functional groups. The knowledge graph comprises 148,507 unique nodes: 147,564 corresponding to ring systems and 943 representing non-ring functional groups. Training is conducted with a batch size of 64, a learning rate of $1\times 10^{-3}$, over 50 epochs, with model checkpoints saved at the end of each epoch.
ComplEx Model Representation
In the ComplEx model [Trouillon et al., 2016], each element in a triple $(\mathbf{h},\mathbf{r},\mathbf{t})$ — where $\mathbf{h}$ is the head entity, $\mathbf{r}$ is the relation, and $\mathbf{t}$ is the tail entity — is represented as a complex vector:
$\mathbf{h},\mathbf{r},\mathbf{t}\in\mathbb{C}^{d}$ (2)
Scoring Function
The score for a given triple $(\mathbf{h},\mathbf{r},\mathbf{t})$ is calculated as:
$f(\mathbf{h},\mathbf{r},\mathbf{t})=\text{Re}\left(\mathbf{h}^{T}\mathbf{r}\cdot\mathbf{t}\right)$ (3)
where $\mathbf{r}$ is a complex-valued vector, and the dot product is performed in the complex space.
Loss Function
ComplEx employs a margin-based ranking loss function defined as:
$\mathcal{L}_{\text{Graph}}=\sum_{(h,r,t)\in E^{+}}\sum_{(\mathbf{h^{\prime}},\mathbf{r},\mathbf{t^{\prime}})\in E^{-}}\max\left(0,\gamma+f(\mathbf{h^{\prime}},\mathbf{r},\mathbf{t^{\prime}})-f(\mathbf{h},\mathbf{r},\mathbf{t})\right)$ (4)
where $E^{+}$ denotes the set of positive triples, $E^{-}$ denotes the set of negative triples, and $\gamma$ represents the margin.
To assess the quality of the learned embeddings, we randomly sample clusters of five closely related embedding vectors and analyze their arrangement in the embedding space. The results of this evaluation are presented in Figure 4a.
D.3 Link Prediction Model Using GNNs
For link prediction using the GCN model, we start by segmenting molecules into functional groups via FG-aware molecular segmentation, where each group is connected by single bonds. We then use embeddings from the FG knowledge graph embedding model as node features for the GCN. The training process involves computing node embeddings through graph convolution (Equation 5), followed by scoring potential edges with a multi-layer perceptron (MLP) (Equation 6). This score is used to calculate the probability between two nodes (Equation 7). Positive and negative edges are sampled, and the model is optimized to maximize scores for positive edges while minimizing scores for negative edges using the loss function in Equation 8. This approach effectively trains the model to distinguish between likely and unlikely connections between functional groups.
$\mathbf{h}_{i}^{\prime}=\text{ReLU}\left(\mathbf{W}\cdot\frac{1}{|\mathcal{N}(i)|}\sum_{j\in\mathcal{N}(i)}\mathbf{h}_{j}\right)$ (5)
where $\mathbf{h}_{i}^{\prime}$ is the updated embedding for node $i$. It is computed by averaging the embeddings $\mathbf{h}_{j}$ of neighboring nodes $\mathcal{N}(i)$, applying the weight matrix $\mathbf{W}$, and then passing through the ReLU activation function.
$s_{ij}=\text{MLP}(\mathbf{h}_{i}\oplus\mathbf{h}_{j})$ (6)
where $s_{ij}$ denotes the score assigned to the potential edge between nodes $i$ and $j$. The score is computed using a multi-layer perceptron (MLP), which takes as input the concatenated node embeddings of $i$ and $j$, denoted as $\mathbf{h}_{i}\oplus\mathbf{h}_{j}$. Here, $\mathbf{h}_{i}$ and $\mathbf{h}_{j}$ represent the node embeddings for nodes $i$ and $j$, respectively. The operator $\oplus$ indicates the concatenation of these embeddings. The MLP processes this concatenated vector to produce a score that reflects the likelihood of an edge existing between $i$ and $j$.
$p_{ij}=\sigma(s_{ij})$ (7)
where $\sigma$ is the sigmoid function.
$\mathcal{L}_{\text{Link}}=-\frac{1}{|E^{+}|}\sum_{(i,j)\in E^{+}}\log p_{ij}-\frac{1}{|E^{-}|}\sum_{(i,j)\in E^{-}}\log(1-p_{ij})$ (8)
where $\mathcal{L}$ is the loss function for link prediction. It computes the average log-likelihood of positive edges $E^{+}$ and negative edges $E^{-}$, where $p_{ij}$ is the predicted probability of an edge between nodes $i$ and $j$. The loss penalizes the model for incorrect predictions, encouraging high probabilities for true edges and low probabilities for false edges.
The GCN model for link prediction is trained as follows: For each molecule, represented as a FG graph, we generate all possible combinations of nodes, encompassing both positive pairs (nodes that are linked) and negative pairs (nodes that are not linked). In cases where the graph contains more than three nodes (FGs), we select 60% of all possible combinations along with all positive pairs to form the training data for each graph. The model is subsequently trained for three epochs on a comprehensive dataset consisting of 20 million data points. Figure 8 shows the performance of the link prediction model. Similar to word embedding analogies in NLP, replacing one FG in a molecule with another produces parallel results across different molecules, demonstrating the model’s ability to capture chemical relationships effectively.
Figure 8: Link prediction model performance: Similar to word embedding analogies in NLP, replacing one functional group in a molecule with another produces parallel results across different molecules, demonstrating the model’s ability to capture chemical relationships effectively.
D.4 Contrastive Learning: Align SMILES and Structure Representation
In this work, we propose a contrastive learning strategy to align SMILES-based representations of molecules with their corresponding graph-based molecular structures. The goal of this approach is to capture both the sequential information from SMILES and the structural relationships encoded in graph representations, thus allowing the model to learn a more comprehensive molecular representation that bridges these two modalities.
To measure the similarity between representations derived from the FG-enhanced SMILES and FG graph, we utilize cosine similarity, which is defined as: The cosine similarity between two vectors $\mathbf{u}$ and $\mathbf{v}$ is defined as:
[…]
Here, $\mathbf{u}$ and $\mathbf{v}$ represent the embeddings from two different modalities, such as the SMILES-based BERT output and the GNN output for the molecular graph. This similarity score helps ensure that embeddings of positive (i.e., matched) SMILES and graph representations are closer in the latent space.
To align these two types of representations, we use contrastive loss, a popular technique in self-supervised learning that enforces representations from the same sample (positive pair) to be more similar than those from different samples (negative pair). Given a positive pair $(\mathbf{h}_{\text{MLM}},\mathbf{h}_{\text{pos}})$, where $\mathbf{h}_{\text{MLM}}$ is the SMILES representation derived from a pretrained BERT model and $\mathbf{h}_{\text{pos}}$ is the corresponding representation from a graph neural network (GNN), and a negative pair $(\mathbf{h}_{\text{MLM}},\mathbf{h}_{\text{neg}})$, where $\mathbf{h}_{\text{neg}}$ is a augmented FG-graph, the contrastive loss can be written as:
$\mathcal{L}_{\text{CL}}=\frac{1}{N}\sum_{i=1}^{N}\max\left(0,\gamma-\text{cos}(\mathbf{h}_{\text{MLM}},\mathbf{h}_{\text{pos}})+\text{cos}(\mathbf{h}_{\text{MLM}},\mathbf{h}_{\text{neg}})\right)$
Where:
  • $\gamma$ is the margin parameter, ensuring that the positive similarity is significantly larger than the negative similarity.
  • $N$ is the number of training examples (or contrastive pairs)
The objective function is
$\mathcal{L}=\lambda_{\text{MLM}}\cdot\mathcal{L}_{\text{MLM}}+\lambda_{\text{CL}}\cdot\mathcal{L}_{\text{CL}}$
where $\mathcal{L}_{\text{MLM}}$ represents the masked language modeling loss, which encourages the model to predict masked tokens in the input sequence effectively, and $\mathcal{L}_{\text{CL}}$ denotes the contrastive loss, which aligns the SMILES and structural representations. The coefficients $\lambda_{\text{MLM}}$ and $\lambda_{\text{CL}}$ are hyperparameters that control the contribution of each loss to the overall objective. By tuning these coefficients, we can balance the learning process between the two tasks, allowing the model to learn rich and meaningful representations from both the sequential and structural aspects of the molecular data.
This combined loss function enables the model to leverage the strengths of both masked language modeling and contrastive learning, fostering a more comprehensive understanding of molecular representations that can enhance performance in downstream tasks such as property prediction, molecular generation, and structure-based drug discovery.
In our contrastive learning model, we set the margin $\gamma=0.5$ and the weights $\lambda_{MLM}=1.0$ and $\lambda_{CL}=0.5$. We train the contrastive BERT model using a batch size of 126 for a total of 5 epochs. This training configuration mirrors the setup used for learning atom representations with the BERT model, as described in Section D.1.
D.5 Downstream task finetuning
MoleculeNet tasks are treated as downstream tasks for our FARM model. We freeze all layers of FARM and pair it with a GRU head for both classification and regression tasks. For classification, we use cross-entropy as the loss function, while for regression, we employ mean squared error. The Adam optimizer is applied with a learning rate of $1e-4$ and a cosine annealing learning rate schedule with a period of 20 epochs. The training process spans 100 epochs with a batch size of 16, using an 80-10-10 train-validation-test split with scaffold splitting. To address imbalanced datasets, we implement a weighted loss function, assigning a weight of 5 to classes with fewer samples. For each task, we conduct three runs with different train-validation-test splits and report the average and standard deviation of the results.
Appendix E Evaluation of FARM on Quantum Mechanical Property Prediction
Table 9: Performance of FARM on quantum mechanical property prediction using the Quantum Mechanical dataset. $R^{2}$ scores are reported with standard deviation over 10 runs.
Class $\mathbf{R^{2}}$ Class $\mathbf{R^{2}}$
HOMOm1 (eV) 0.927±0.01480.927$\pm$ 0.01480.927 ± 0.0148 T5 0.905±0.03150.905$\pm$ 0.03150.905 ± 0.0315
HOMO (eV) 0.895±0.02980.895$\pm$ 0.02980.895 ± 0.0298 S1 0.936±0.01610.936$\pm$ 0.01610.936 ± 0.0161
LUMO (eV) 0.928±0.009480.928$\pm$ 0.009480.928 ± 0.00948 S2 0.948±0.01560.948$\pm$ 0.01560.948 ± 0.0156
LUMOp1 (eV) 0.922±0.02800.922$\pm$ 0.02800.922 ± 0.0280 S3 0.911±0.008490.911$\pm$ 0.008490.911 ± 0.00849
DipoleMoment (Debye) 0.755±0.05320.755$\pm$ 0.05320.755 ± 0.0532 S4 0.919±0.007920.919$\pm$ 0.007920.919 ± 0.00792
LogP 0.995±0.005660.995$\pm$ 0.005660.995 ± 0.00566 S5 0.924±0.01430.924$\pm$ 0.01430.924 ± 0.0143
PrimeExcite (eV) 0.877±0.04910.877$\pm$ 0.04910.877 ± 0.0491 O1 0.839±0.03400.839$\pm$ 0.03400.839 ± 0.0340
PrimeExcite (osc) 0.837±0.02430.837$\pm$ 0.02430.837 ± 0.0243 TDOS2.0 0.868±0.09900.868$\pm$ 0.09900.868 ± 0.0990
T1 0.929±0.005870.929$\pm$ 0.005870.929 ± 0.00587 TDOS2.5 0.877±0.004100.877$\pm$ 0.004100.877 ± 0.00410
T2 0.925±0.01820.925$\pm$ 0.01820.925 ± 0.0182 TDOS3.0 0.815±0.04470.815$\pm$ 0.04470.815 ± 0.0447
T3 0.900±0.01010.900$\pm$ $0$ 0.0101 TDOS3.5 0.895±0.03270.895$\pm$ 0.03270.895 ± 0.0327
T4 0.921±0.01960.921$\pm$ 0.01960.921 ± 0.0196 TDOS4.0 0.919±0.03130.919$\pm$ 0.03130.919 ± 0.0313
TDOS4.5 0.938±0.01830.938$\pm$ 0.01830.938 ± 0.0183
Table 9 presents the results of FARM on the quantum mechanical property prediction tasks from the dataset introduced in [Angello et al., 2024]. FARM demonstrates consistently high performance across a wide range of properties, including difficult-to-predict targets such as Dipole Moment, excited-state energies (S1–S5 and T1–T5), and total density of states (TDOS). These results underscore FARM’s ability to accurately model complex quantum mechanical behaviors by leveraging functional group information. Its success in capturing nuanced molecular interactions further illustrates FARM’s applicability to real-world challenges in chemical property prediction, where understanding the role of functional modules is essential for robust and generalizable predictions.
Appendix F Ablation Study
F.1 FARM Component Analysis
To assess the effectiveness of each component in our architecture, we conducted a comprehensive ablation study across several MoleculeNet benchmark tasks. The first model, FM_KGE + GAT, utilizes FG knowledge graph embeddings as input for a Graph Attention Network [Veličković et al., 2017] (GAT) to predict molecular properties. Although its performance on these tasks is not the strongest, the model still demonstrates its capacity to learn underlying chemical rules (syntax and semantics) from the data to a certain degree.
The second model, AttentiveFP [Xiong et al., 2019], performs a masked atom prediction task on the molecular graph, predicting atom types such as carbon, hydrogen, oxygen, and nitrogen. Its variation, FG AttentiveFP, shares the same architecture as AttentiveFP, but it predicts both the atom type and the associated functional group. Experimental results indicate that incorporating functional group information significantly improves the model’s performance on downstream tasks.
We also evaluate the BERT model trained on canonical SMILES strings, and its counterpart, FG BERT, which is trained on FG-enhanced SMILES. Results show that providing additional chemical context about functional groups boosts model performance in downstream tasks.
Finally, FARM (FG BERT with contrastive learning) integrates molecular structure representations from link prediction embeddings. FARM consistently achieves the highest performance across 6 out of 7 downstream tasks, demonstrating the power of combining FG-enhanced SMILES and contrastive learning.
Table 10 presents the detailed results of the aforementioned models across various MoleculeNet tasks, illustrating the performance of each architecture. For these experiments, we used random splitting to divide the downstream datasets into training, validation, and test sets in an 8:1:1 ratio. While random splitting is used consistently across models in this ablation study, scaffold splitting is applied for benchmarking to ensure a fair comparison with other methods.
Table 10: Performance of various models across 6 MoleculeNet tasks. The data is split using a random split into training, validation, and test sets with an 8:1:1 ratio.
BBBP BACE HIV Average ESOL FreeSolv Average QM9
#tasks 1 1 1 1 1 3
#samples 2039 1513 41127 1128 642 133885
Metric ROC-AUC ($\uparrow$) RMSE ($\downarrow$) MAE ($\downarrow$)
FG_KGE + GAT 73.23 $\pm$ 1.93 76.44 $\pm$ 1.27 71.65 $\pm$ 0.98 73.77 2.35 $\pm$ 0.210 4.32 $\pm$ 0.29 3.335 0.0139 $\pm$ 0.00014
AttentiveFP 77.71 $\pm$ 1.30 77.15 $\pm$ 0.78 78.81 $\pm$ 0.99 77.89 1.63 $\pm$ 0.042 2.11 $\pm$ 0.94 1.87 0.0056 $\pm$ 0.00012
FG AttentiveFP 85.57 $\pm$ 1.32 87.30 $\pm$ 0.90 81.21 $\pm$ 0.92 84.5 1.02 $\pm$ 0.034 1.08 $\pm$ 0.14 1.05 0.0053 $\pm$ 0.00034
BERT 82.12 $\pm$ 1.45 85.12 $\pm$ 0.76 83.03 $\pm$ 1.12 83.42 1.45 $\pm$ 0.056 1.89 $\pm$ 0.09 1.67 0.0059 $\pm$ 0.00012
FG BERT 94.36 $\pm$ 0.50 94.54 $\pm$ 0.40 81.93 $\pm$ 1.70 90.27 0.608 $\pm$ 0.031 0.507 $\pm$ 0.03 0.558 0.0041 $\pm$ 0.00017
FARM 96.23 $\pm$ 0.7 96.19 $\pm$ 0.65 82.13 $\pm$ 1.10 91.43 0.734 $\pm$ 0.039 0.308 $\pm$ 0.08 0.521 0.0038 $\pm$ 0.00014
F.2 Fragmentation Methods
In this section, we compare the performance of our FG-based fragmentation method with BRICS fragmentation. When using BRICS as the fragmentation method and enriching SMILES with functional group information derived from BRICS (treating each BRICS fragment as a functional group), the resulting vocabulary size is 575,612 (with a minimum frequency of 5), compare to 14,741 with our FG-based fragmentation. This significant difference arises because BRICS tends to generate larger fragments, causing the vocabulary to grow rapidly. This larger vocabulary can make the model harder to train and converge, as it includes a high number of rare tokens. Consequently, the likelihood of encountering out-of-vocabulary tokens during validation and testing increases, which negatively impacts model performance. As shown in Table 4, under the same experimental setup, our FG-based fragmentation method outperforms BRICS fragmentation across all six MoleculeNet tasks.
Appendix G Computational Resources
For the BERT model, even with an expanded vocabulary of approximately 14,000 tokens, the model size remains smaller than the English BERT vocabulary and does not require additional computational resources compared to training a standard BERT model. Each BERT model, with around 100 million parameters, was trained for 20 epochs on a machine with an NVIDIA A100 GPU (40GB memory), 256GB of RAM, and a 32-core AMD EPYC 7502 CPU. Training each model took approximately 6 hours. The experiments were conducted on an internal computing cluster, and no additional compute was required beyond the primary experiments included in the paper.
For other models, such as those used for ComplEx and link prediction GNN, the computational requirements are relatively lightweight, and the same machine used for BERT training was sufficient without the need for additional resources.